### v1

@@
The existence of extremely massive and compact galaxies (Little Red Dots or LRDs) at $z\gtrsim 2$ challenges models of early structure formation, suggesting extremely rapid stellar and black hole (BH) assembly. These galaxies are efficient environments for rapid BH growth, yet several LRDs show no evidence of strong emission from the active galactic nuclei (AGN) in X-rays. Our work utilizes a subsample of X-ray non-detected LRDs to check whether the collision-based BH formation scenario is compatible with the non-detections and constrain both physical parameters (e.g., metallicity) and observational parameters (e.g., column density). Our results show that LRDs might be ideal birthplaces to form massive BHs, particularly in the case of a mass-radius relation $R_{\rm gal}\propto M^{0.6}_{\rm gal}$ (similar to spiral galaxies in the Local Universe). Given the large stellar densities, collision-based models suggest seed masses greater than those observed in the local Universe and those are compatible with the mass-radius relation of high-redshift BHs. We model the BH seed formation and subsequent X-ray emission investigating both the physical and observational parameter space to compare with the observed X-ray upper limits in the soft band (0.320.3-2 keV) and the hard band (272-7 keV). We find exponents in the mass-radius relation above $0.55$ favor the collision-based scenario; however, consistency with the stacked X-ray analysis requires specific combinations of accretion and obscuration parameters. Additionally, we find that constant or increasing star formation rate scenarios assuming high Eddington ratios and large duty cycles are feasible but require larger column densities and/or a higher level of metal enrichment in the shielding columns. Alternatively, moderate sub-Eddington accretion rates seem to be sufficient to reconcile the massive seeds with their final observed masses, consistent with the observed X-ray weakness. Overall, we conclude that even in cases where LRDs were initially star-burst galaxies, they should evolve into an AGN.
@@ 1 Introduction
The detection of extremely compact and massive galaxies at high redshift by JWST challenges the current knowledge of galaxy formation and evolution. These galaxies, the so-called Little Red Dots (LRDs), have masses ranging between $10^{8}$ and $10^{12}$ M in an effective radius of 33003-300 pc (e.g., Matthee et al., 2024; Greene et al., 2024; Akins et al., 2024). What characterizes an LRD is its ‘V-shaped’ Spectral Energy Distribution (SED), which features an inclined red continuum in the optical rest frame together with a blue UV counterpart in the rest frame (Kocevski et al., 2023; Akins et al., 2023; Maiolino et al., 2024). The large sensitivity of JWST surveys, particularly the JWST Advanced Deep Extragalactic Survey (JADES) (Eisenstein et al., 2023a, b), facilitates the discovery and characterization of LRDs over the GOODS-South and GOODS-North fields, contributing to the measurement of their properties (e.g., Rinaldi et al., 2025). However, the origin of this peculiar spectrum is still strongly debated. The slopes in the optical band are consistent with (i) emission from dusty star-formation (i.e., emission dominated by young stellar populations; Williams et al., 2024; Pérez-González et al., 2024) or (ii) reddened Active Galactic Nuclei (AGN) where thermal emission from the accretion disk dominates on scales less than $1$ pc with temperatures higher than $10^{5}$ K (Labbe et al., 2025; Matthee et al., 2024). Similarly, the excess of UV emission might be explained by light from the central AGN and stellar emission emerging from a relatively dust-free host, or by light escaping without attenuation due to patchy dust, as observed in some lower redshift dusty star-forming galaxies (e.g., Casey et al., 2014) and red quasars (Glikman et al., 2023).
@@
While the broad emission lines (FWHM $\sim 2000\,{\rm km\,s}^{-1}$) observed in many LRDs suggest the presence of accreting supermassive black holes (SMBHs) with masses of at least $10^{7}-10^{8}$ M (Matthee et al., 2024), confirming this AGN nature in the X-ray regime has proven difficult. Recent stacking analyses of X-ray non-detected sources have yielded strict upper limits. For instance, Sacchi and Bogdan (2025) (hereafter S+25) constrained the emission of galaxies lacking individual X-ray detections. They found upper limits that are incompatible with unobscured super-Eddington accretion models. Their results imply that if massive BHs are present, they must be either significantly less massive than optical estimates suggest, or are buried under extreme column densities ($N_{\rm H}\gtrsim 10^{25}$ cm-2). This characteristic “X-ray weakness” is supported by independent studies. Ananna et al. (2024) found no significant detection in a stack of 21 high-redshift LRDs, placing upper limits on BH masses of $\lesssim(1.5-16)\times 10^{6}$ M assuming Eddington-limited accretion. Similarly, Yue et al. (2024) reported only tentative detections in a stack of 34 LRDs, suggesting that any central engine must be intrinsically weak or heavily obscured. This discrepancy between the optical evidence for massive BHs and the general lack of strong X-ray emission presents a major open question regarding the nature and growth of LRDs.
@@ 2.1 Population data
10* $0.46^{+0.05}_{-0.05}$ $1.023^{+0.2}_{-0.2}$ 0.30 dex Stefanon et al. (2021)
Table 1: Summary of the parameters used in the scaling laws (Eq. 1) to estimate the stellar mass of the galaxies. (1) Redshift range $z$; (2) slope of the correlation $a$; (3) intercept coefficient $b$; (4) intrinsic scatter; (5) Reference for the values adopted here. (*) As there is no published correlation for redshift 11 we use the same as for $z=10$ in Stefanon et al. (2021).
As the radii for the sources found in S+25 sample have not been yet published, we assume here that they follow the same distribution of the sample in A+24, where it appears roughly independent of mass (see Appendix A). We obtain the effective radius of the galaxies in A+24 sample. In there, the authors multiply the radius (in mas) by the angular distance $D_{A}(z)$. As galaxy sizes span on several orders of magnitude, we performed the fitting in logarithmic space, defined as $x=\log_{10}(R_{\text{gal}}/\text{pc})$. We employed a Gaussian Mixture Model approach, where the distribution is described as the sum of $K$ independent Gaussian components. The probability density $P(x)$ is given by $P(x)=\sum_{k=1}^{K}w_{k}\frac{1}{\sqrt{2\pi\sigma_{k}^{2}}}\exp\left(-\frac{(x-\mu_{k})^{2}}{2\sigma_{k}^{2}}\right),$ where $w_{k}$, $\mu_{k}$, and $\sigma_{k}$ are the weight, mean, and standard deviation of the $k$-th component, respectively, subject to the constraint $\sum w_{k}=1$. The parameters were estimated using the Expectation-Maximization algorithm. We compared models with $K=1$ and $K=2$ components (see Fig. 2), evaluating the goodness-of-fit using the reduced chi-squared statistic ($\chi^{2}_{\nu}$) calculated over the histogram bins assuming a fixed size equals to $0.2$  dex. Prior to fitting, we excluded a single object with an effective radius $R_{\text{eff}}>1$ kpc. As this object was the sole galaxy in the sample with such a large extent, treating it as part of the continuous distribution would disproportionately affect the parametric fit or require an statistically unjustified third component; removing it allows for a robust characterization of the primary compact and extended populations. Even though the redshift and the mass distributions of the A+24 sample and the S+25 sub-sample are slightly different, we assume that the radius distribution used here provides a conservative estimate as discussed in Appendix B. We checked how results are affected by sampling from the one and two components distributions finding no relevant changes.
@@ 2.2 Dynamics of Little Red Dots
where $\beta$ is the exponent of the correlation $R_{\rm gal}-M_{\rm gal}$ [^1] Note that the normalization of the correlation is individual for each source in order to match the final radius sampled from the distribution. and assumed to be in the range of 0.11.50.1-1.5. This parameter is a key point of observational debate at high redshift. For example, Allen et al. (2025) measured a constant size-mass relation slope of $\beta\approx 0.215$ for star-forming galaxies from z=39z=3-9. In contrast, Ormerod et al. (2024) found that this relationship breaks down at $z>3$, with galaxy sizes showing little to no dependence on stellar mass (implying $\beta\approx 0$). For NSCs, we have $\beta\sim 0.5$, while early-type galaxies have typical values of $\beta=0.55-0.6$ (Shen et al., 2003) and late-type galaxies have $\beta=0.22-0.27$ (Shen et al., 2003; Lange et al., 2015). Our model, which tests a wide range of $\beta$, can therefore provide constraints on which evolutionary track is more consistent with the observed properties of LRDs.
@@ 2.4 X-ray background data and black hole emission
We model the intrinsic X-ray emission of the SMBHs formed in our model. We assume a power law with a high-energy exponential cut-off (e.g., Yang et al., 2020)
@@
where $\Gamma$ is the photon index. The value of $\Gamma$ is usually set at $\approx 1.8-1.9$ (e.g., Piconcelli et al., 2005; Yang et al., 2016; Liu et al., 2017; Yang et al., 2020) and $E_{\rm cut}=300$ keV, for Seyfert galaxies (see Ricci et al., 2017). However, the weak emission of LRDs in the X-ray band is consistent with a steeper value of $\Gamma\approx 2.4\pm 0.1$ (e.g., Zappacosta et al., 2023) or with the presence of a very low energy cutoff $E_{\rm cut}\approx 20$  keV. We adopt the canonical value $\Gamma=1.9$ with a low energy cut-off of $E_{\rm cut}=20$  keV. We also checked the possibility of $\Gamma=2.4$ finding no difference in our results.
@@ 3.2.2 Metallicity and average column density
In Fig. 8, we present the results for a value of $k=0.3$, corresponding to a case with high Eddington ratio, duty cycle and BH population fraction in the top panel, and we show the results of a value equals to $k=2\times 10^{-3}$ for a more conservative scenario (bottom panel). The colorbar indicates the ratio of predicted versus observed flux in the soft band. So far, no metallicities of LRDs have been inferred from observations. Furthermore, observations at low and high redshift suggest that the range of metallicities is not constrained by redshift. Indeed, at high redshift, highly evolved objects like AGNs also show close-to-solar metallicities (e.g., Nagao et al., 2006).
@@ 3.2.4 Black hole activity and metallicity
On the other hand, for high redshift BHs we may expect larger values of the duty cycle and the Eddington ratio, even if the precise value of $k$ is unconstrained. The situation improves quite relevantly with a larger column density of $N_{\rm ave}=10^{21}$ cm-2, for which even values of $k\sim 0.3$ become compatible with the constraints in case that the metallicity is high. Even at lower metallicities this effect can likely be mitigated by larger gas columns increasing the amount of shielding. The overall constraints for this case are summarized in Fig. 10 as a function of metallicity and as a function of $k$.
@@ Appendix A MCMC fitting
Table 2: The uniform priors $\pi(\theta)$ adopted for the model parameters $\theta=(A,B,\log_{10}\sigma)$.
@@
Table 3: Best fit parameters from our MCMC analysis. We show the results of the best values plus the $1\sigma$ errors.
@@ Appendix B Samples comparison
In this Appendix, we compare the properties of the A+24 sample, which we use to define the structural parameters of Little Red Dots (LRDs), and the sub-sample of S+25 used for the X-ray analysis in this work.
In Fig. 15, we compare the stellar mass distributions of the two samples. The A+24 sample (red) is peaked at high masses ($M_{\rm gal}\sim 10^{10}-10^{11}$ M), while the S+25 sub-sample (blue) exhibits an extended tail of low-mass objects reaching down to $10^{6}$ M. We argue that the inclusion of lower-mass galaxies in the S+25 sample does not introduce a bias in our modeling. This is due to the inherent nature of X-ray stacking, where the aggregate signal is flux-weighted. Since the X-ray luminosity scales with stellar mass ($L_{X}\propto M_{\rm gal}$, e.g., Anderson et al. 2015), the stacked signal is dominated by the most massive galaxies in the sample. A single galaxy with a mass of $10^{10}$ M contributes a flux comparable to thousands of $10^{6}$ M systems. Therefore, the structural parameters of the massive end of the population (represented by the A+24 sample) are the relevant priors for modeling the sources responsible for the potential X-ray emission.
Figure 15: Comparison of the stellar mass distributions between the A+24 sample (red) and the S+25 sub-sample (blue). While the S+25 sample extends to lower masses, the X-ray stacking signal is dominated by the massive end of the population, justifying the use of structural parameters derived from the A+24 sample.

### v2

@@
The existence of extremely massive and compact galaxies (Little Red Dots or LRDs) at $z\gtrsim 2$ challenges models of early structure formation, suggesting extremely rapid stellar and black hole (BH) assembly. These galaxies are efficient environments for rapid BH growth, yet several LRDs show no evidence of strong emission from the active galactic nuclei (AGN) in X-rays. Our work utilizes a subsample of X-ray non-detected LRDs to check whether the collision-based BH formation scenario is compatible with the non-detections and constrain both physical parameters (e.g., metallicity) and observational parameters (e.g., column density). Our results show that LRDs might be ideal birthplaces to form massive BHs, particularly in the case of a mass-radius relation $R_{\rm gal}\propto M^{0.6}_{\rm gal}$ (similar to spiral galaxies in the Local Universe). Given the large stellar densities, collision-based models suggest seed masses greater than those observed in the local Universe and those are compatible with the mass-radius relation of high-redshift BHs. We model the BH seed formation and subsequent X-ray emission investigating both the physical and observational parameter space to compare with the observed X-ray upper limits in the soft band (0.320.3-2 keV) and the hard band (272-7 keV). We find exponents in the mass-radius relation above $0.55$ favor the collision-based scenario; however, consistency with the stacked X-ray analysis requires specific combinations of accretion and obscuration parameters. Additionally, we find that constant or increasing star formation rate scenarios assuming high Eddington ratios and large duty cycles are feasible but require larger column densities and/or a higher level of metal enrichment in the shielding columns. Alternatively, moderate sub-Eddington accretion rates seem to be sufficient to reconcile the massive seeds with their final observed masses, consistent with the observed X-ray weakness. Overall, we conclude that even in cases where LRDs were initially starburst galaxies, they should evolve into an AGN.
@@ 1 Introduction
The detection of extremely compact and massive galaxies at high-redshift by JWST challenges the current knowledge of galaxy formation and evolution. These galaxies, the so-called Little Red Dots (LRDs), have masses ranging between $10^{8}$ and $10^{12}$ M in an effective radius of 33003-300 pc (e.g., Matthee et al., 2024; Greene et al., 2024; Akins et al., 2025). What characterizes an LRD is its ‘V-shaped’ Spectral Energy Distribution (SED), which features an inclined red continuum in the optical rest frame together with a blue UV counterpart in the rest frame (Kocevski et al., 2023; Akins et al., 2023; Maiolino et al., 2024). The large sensitivity of JWST surveys, particularly the JWST Advanced Deep Extragalactic Survey (JADES) (Eisenstein et al., 2025, 2026), facilitates the discovery and characterization of LRDs over the GOODS-South and GOODS-North fields, contributing to the measurement of their properties (e.g., Rinaldi et al., 2025). However, the origin of this peculiar spectrum is still strongly debated. The slopes in the optical band are consistent with (i) emission from dusty star-formation (i.e., emission dominated by young stellar populations; Williams et al., 2024; Pérez-González et al., 2024) or (ii) reddened Active Galactic Nuclei (AGN) where thermal emission from the accretion disk dominates on scales less than $1$ pc with temperatures higher than $10^{5}$ K (Labbe et al., 2025; Matthee et al., 2024). Similarly, the excess of UV emission might be explained by light from the central AGN and stellar emission emerging from a relatively dust-free host, or by light escaping without attenuation due to patchy dust, as observed in some lower redshift dusty star-forming galaxies (e.g., Casey et al., 2014) and red quasars (Glikman et al., 2023).
@@
While the broad emission lines (FWHM $\sim 2000\,{\rm km\,s}^{-1}$) observed in many LRDs suggest the presence of accreting supermassive black holes (SMBHs) with masses of at least $10^{7}-10^{8}$ M (Matthee et al., 2024), confirming this AGN nature in the X-ray regime has proven difficult. Recent stacking analyses of X-ray non-detected sources have yielded strict upper limits. For instance, Sacchi and Bogdán (2025) (hereafter S+25) constrained the emission of galaxies lacking individual X-ray detections. They found upper limits that are incompatible with unobscured super-Eddington accretion models. Their results imply that if massive BHs are present, they must be either significantly less massive than optical estimates suggest, or are buried under extreme column densities ($N_{\rm H}\gtrsim 10^{25}$ cm-2). This characteristic “X-ray weakness” is supported by independent studies. Ananna et al. (2024) found no significant detection in a stack of 21 high-redshift LRDs, placing upper limits on BH masses of $\lesssim(1.5-16)\times 10^{6}$ M assuming Eddington-limited accretion. Similarly, Yue et al. (2024) reported only tentative detections in a stack of 34 LRDs, suggesting that any central engine must be intrinsically weak or heavily obscured. This discrepancy between the optical evidence for massive BHs and the general lack of strong X-ray emission presents a major open question regarding the nature and growth of LRDs.
@@ 2.1 Population data
Table 1: Summary of the parameters used in the scaling laws (Eq. 1) to estimate the stellar mass of the galaxies.111(1) Redshift range $z$; (2) slope of the correlation $a$; (3) intercept coefficient $b$; (4) intrinsic scatter; (5) Reference for the values adopted here.
@@
10aaaaaaAs there is no published correlation for redshift 11 we use the same as for $z=10$ in Stefanon et al. (2021). $0.46^{+0.05}_{-0.05}$ $1.023^{+0.2}_{-0.2}$ 0.30 dex Stefanon et al. (2021)
As the radii for the sources found in S+25 sample have not been yet published, we assume here that they follow the same distribution of the sample in A+24, where it appears roughly independent of mass (see Appendix A). We obtain the effective radius of the galaxies in A+24 sample. In there, the authors multiply the radius (in mas) by the angular distance $D_{A}(z)$. As galaxy sizes span on several orders of magnitude, we performed the fitting in logarithmic space, defined as $x=\log_{10}(R_{\text{gal}}/\text{pc})$. We employed a Gaussian Mixture Model approach, where the distribution is described as the sum of $K$ independent Gaussian components. The probability density $P(x)$ is given by $P(x)=\sum_{k=1}^{K}w_{k}\frac{1}{\sqrt{2\pi\sigma_{k}^{2}}}\exp\left(-\frac{(x-\mu_{k})^{2}}{2\sigma_{k}^{2}}\right),$ where $w_{k}$, $\mu_{k}$, and $\sigma_{k}$ are the weight, mean, and standard deviation of the $k$-th component, respectively, subject to the constraint $\sum w_{k}=1$. The parameters were estimated using the Expectation-Maximization algorithm. We compared models with $K=1$ and $K=2$ components (see Fig. 2), evaluating the goodness-of-fit using the reduced chi-squared statistic ($\chi^{2}_{\nu}$) calculated over the histogram bins assuming a fixed size equals to $0.2$  dex. Prior to fitting, we excluded a single object with an effective radius $R_{\text{eff}}>1$ kpc. As this object was the sole galaxy in the sample with such a large extent, treating it as part of the continuous distribution would disproportionately affect the parametric fit or require an statistically unjustified third component; removing it allows for a robust characterization of the primary compact and extended populations. Even though the redshift and the mass distributions of the A+24 sample and the S+25 subsample are slightly different, we assume that the radius distribution used here provides a conservative estimate as discussed in Appendix B. We checked how results are affected by sampling from the one and two components distributions finding no relevant changes.
@@ 2.2 Dynamics of Little Red Dots
where $\beta$ is the exponent of the correlation $R_{\rm gal}-M_{\rm gal}$ [^2] Note that the normalization of the correlation is individual for each source in order to match the final radius sampled from the distribution. and assumed to be in the range of 0.11.50.1-1.5. This parameter is a key point of observational debate at high-redshift. For example, Allen et al. (2025) measured a constant size-mass relation slope of $\beta\approx 0.215$ for star-forming galaxies from z=39z=3-9. In contrast, Ormerod et al. (2024) found that this relationship breaks down at $z>3$, with galaxy sizes showing little to no dependence on stellar mass (implying $\beta\approx 0$). For NSCs, we have $\beta\sim 0.5$, while early-type galaxies have typical values of $\beta=0.55-0.6$ (Shen et al., 2003) and late-type galaxies have $\beta=0.22-0.27$ (Shen et al., 2003; Lange et al., 2015). Our model, which tests a wide range of $\beta$, can therefore provide constraints on which evolutionary track is more consistent with the observed properties of LRDs.
@@ 2.4 X-ray background data and black hole emission
We model the intrinsic X-ray emission of the SMBHs formed in our model. We assume a power law with a high-energy exponential cutoff (e.g., Yang et al., 2020)
@@
where $\Gamma$ is the photon index. The value of $\Gamma$ is usually set at $\approx 1.8-1.9$ (e.g., Piconcelli et al., 2005; Yang et al., 2016; Liu et al., 2017; Yang et al., 2020) and $E_{\rm cut}=300$ keV, for Seyfert galaxies (see Ricci et al., 2017). However, the weak emission of LRDs in the X-ray band is consistent with a steeper value of $\Gamma\approx 2.4\pm 0.1$ (e.g., Zappacosta et al., 2023) or with the presence of a very low-energy cutoff $E_{\rm cut}\approx 20$  keV. We adopt the canonical value $\Gamma=1.9$ with a low-energy cutoff of $E_{\rm cut}=20$  keV. We also checked the possibility of $\Gamma=2.4$ finding no difference in our results.
@@ 3.2.2 Metallicity and average column density
In Fig. 8, we present the results for a value of $k=0.3$, corresponding to a case with high Eddington ratio, duty cycle and BH population fraction in the top panel, and we show the results of a value equals to $k=2\times 10^{-3}$ for a more conservative scenario (bottom panel). The colorbar indicates the ratio of predicted versus observed flux in the soft band. So far, no metallicities of LRDs have been inferred from observations. Furthermore, observations at low and high-redshift suggest that the range of metallicities is not constrained by redshift. Indeed, at high-redshift, highly evolved objects like AGNs also show close-to-solar metallicities (e.g., Nagao et al., 2006).
@@ 3.2.4 Black hole activity and metallicity
On the other hand, for high-redshift BHs we may expect larger values of the duty cycle and the Eddington ratio, even if the precise value of $k$ is unconstrained. The situation improves quite relevantly with a larger column density of $N_{\rm ave}=10^{21}$ cm-2, for which even values of $k\sim 0.3$ become compatible with the constraints in case that the metallicity is high. Even at lower metallicities this effect can likely be mitigated by larger gas columns increasing the amount of shielding. The overall constraints for this case are summarized in Fig. 10 as a function of metallicity and as a function of $k$.
@@ Appendix A MCMC fitting
Table 2: The uniform priors $\pi(\theta)$ adopted for the MCMC analysis. [^3] The model parameters are $\theta=(A,B,\log_{10}\sigma)$.
@@
Table 3: Best fit parameters from MCMC analysis. [^4] We show the results of the best values plus the $1\sigma$ errors.
@@ Appendix B Samples comparison
In this Appendix, we compare the properties of the A+24 sample, which we use to define the structural parameters of Little Red Dots (LRDs), and the subsample of S+25 used for the X-ray analysis in this work.
In Fig. 15, we compare the stellar mass distributions of the two samples. The A+24 sample (red) is peaked at high masses ($M_{\rm gal}\sim 10^{10}-10^{11}$ M), while the S+25 subsample (blue) exhibits an extended tail of low-mass objects reaching down to $10^{6}$ M. We argue that the inclusion of lower-mass galaxies in the S+25 sample does not introduce a bias in our modeling. This is due to the inherent nature of X-ray stacking, where the aggregate signal is flux-weighted. Since the X-ray luminosity scales with stellar mass ($L_{X}\propto M_{\rm gal}$, e.g., Anderson et al. 2015), the stacked signal is dominated by the most massive galaxies in the sample. A single galaxy with a mass of $10^{10}$ M contributes a flux comparable to thousands of $10^{6}$ M systems. Therefore, the structural parameters of the massive end of the population (represented by the A+24 sample) are the relevant priors for modeling the sources responsible for the potential X-ray emission.
Figure 15: Comparison of the stellar mass distributions between the A+24 sample (red) and the S+25 subsample (blue). While the S+25 sample extends to lower masses, the X-ray stacking signal is dominated by the massive end of the population, justifying the use of structural parameters derived from the A+24 sample.